Find the equation of the line that goes through the given point
and has the given slope. Write the answer in slope-intercept
form.
(-1, -5), -8

thank you

Answer by stanbon(75887)

Find the equation of the line that goes through the given point
and has the given slope. Write the answer in slope-intercept
form.
(-1, -5), -8
---------------------
Form: y = mx + b
-5 = -8(-1)+b
-5 = 8 + b
b = -13
-------
Equation:
y = -8x - 13
==================
Cheers,
Stan H.
=============